OpenStudy (anonymous):

Taylor series for f(x)=ln(sec(x)) at a=0

OpenStudy (abb0t):

take a few derivatives. to find a general formula.

OpenStudy (abb0t):

the general taylor for natural log should be in your book! you can follow that to get a general formula at a=0 after finding a few derivatives and plugging in 0

OpenStudy (anonymous):

right o, i have done that i found f(x)=ln(sec(x), f'(x)=tan(x), f''(x)=sec^2(x), f'''(x)=2tan(x)sec^2(x), f''''(x)=-2(cos(2x)-2)sec^4(x)

OpenStudy (anonymous):

@phi @satellite73 could you try and take a look when you get a chance, please?

OpenStudy (anonymous):

I am also having issues using my ti-83 to calculate either Cn or Sn if you can help with getting the terms to pull up in a table or just by calculating

OpenStudy (anonymous):

i have an idea

OpenStudy (anonymous):

because doing this by hand is really going to suck instead, take the derivative, and you get something nice, namely \[\frac{d}{dx}\ln(\sec(x))=\tan(x)\]

OpenStudy (anonymous):

now the power series for \(\tan(x)\) is well known take that one, and integrate it term by term it will be much much easier this way i think

OpenStudy (anonymous):

that will help, let me give it a shot. would you happen to know how to plug in f(x) into calc ti-83 to get a table showing either series or sum of series?

OpenStudy (anonymous):

no sorry but if you are just looking for a calculator answer (calculators are so last century) you can do this http://www.wolframalpha.com/input/?i=ln%28sec%28x%29%29

OpenStudy (anonymous):

well she also wants us to know how to calculate the error Find the exact error in approximating ln(sec(−0.3)) by its fourth degree Taylor polynomial at a=0